# maximum a posteriori

### v1

@@ D.5 Other gradient penalties and related methods
  • One‑sided GP (Lipschitz penalty). Penalize only $\|\nabla\tilde{h}_{\theta}\|_{2}>1$: $\lambda\,\mathbb{E}[\max(0,\|\nabla\tilde{h}_{\theta}\|-1)^{2}]$, avoiding a push toward <1<1 gradients that can reduce capacity (Petzka et al.,, 2018).
  • Zero‑centered penalties $R_{1}/R_{2}$. Apply $\lambda\,\mathbb{E}_{x\sim P_{\mathrm{data}}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{1}$) or $\lambda\,\mathbb{E}_{x\sim P_{g}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{2}$) to improve convergence empirically—even though they do not strictly enforce $1$-Lipschitzness (Mescheder et al.,, 2018).
  • DRAGAN (local penalties). Sample $\hat{x}=x_{\mathrm{real}}+\delta$ with small noise and penalize $\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}$ near the data manifold to promote local smoothness (Kodali et al.,, 2017).
  • Orthogonal/Parseval constraints. Constrain $W^{\top}W\approx I$ (Parseval/Björck) to shrink $\|W\|_{2}$ toward $1$ at higher compute cost (Cissé et al.,, 2017).
  • Lipschitz activations. Using activations with known Lipschitz constants (e.g., GroupSort) can yield provably $1$-Lipschitz networks when combined with spectral control of linear layers (Anil et al.,, 2019).
@@ Appendix E Additional Experimental Results
E.1 CelebA–64 $\lambda_{\mathrm{GP}}$ Sensitivity for BOLT–GAN
For completeness, Figure 4 presents the CelebA–64 $\lambda_{\mathrm{GP}}$ sweep mirroring the CIFAR-10 analysis in the main text. The stability/performance trends are consistent: $\lambda_{\mathrm{GP}}\!\in\![5,10]$ is best, with variations within $\pm 1$ FID.
@@ E.2 Qualitative Results
We provide qualitative samples to complement the quantitative results in the main paper: each panel below shows an uncurated $4\times 4$ grid of random images generated by BOLT–GAN–GP after 20 training epochs under the same setup as our primary experiments (same architectures, optimizer, and $\lambda_{\mathrm{GP}}$).
Figure 5: Qualitative samples (first experiment of the main paper). BOLT-GAN-GP after 20 epochs with $\lambda_{\mathrm{GP}}{=}10$. Top-left: CIFAR-10; top-right: CelebA-64; bottom-left: LSUN Bedroom-64; bottom-right: LSUN Church-64. Each panel shows the same $4\times 4$ grid of generated samples at a reduced scale. Under the identical setup used in our main experiments, Lipschitz BOLT–GAN produces sharper textures, fewer checkerboard artifacts, and more stable color statistics than the WGAN–GP baseline.
@@ F.4 Objectives and Priors
For reference, we also evaluate a non‑Lipschitz baseline by setting the gradient‑penalty coefficient to zero $(\lambda_{\mathrm{GP}}=0)$. As expected, this variant is numerically unstable and yields high FID; see Sec. S5 for details.
@@ F.5 Lipschitz Enforcement
to the critic objective. We set $\lambda_{\mathrm{GP}}=10$ as our default because this value is the de facto choice introduced by WGAN‑GP and widely adopted in subsequent implementations; it offers a robust trade‑off between stability and capacity in practice (Gulrajani et al.,, 2017; Mescheder et al.,, 2018). In Sec. S5 we also report a small sweep over $\lambda_{\mathrm{GP}}\in\{1,5,10,20\}$ that confirms the default is near‑optimal under our settings.
@@ F.8 Hyperparameter Sweeps and Selection
We sweep $\lambda_{\mathrm{GP}}\in\{1,5,10,20\}$ separately for each method and dataset, and we report the best FID per method within a fixed compute budget. Other knobs such as $\beta_{1}$ and $n_{\mathrm{d}}$ have a comparatively minor effect in our ranges (e.g., within $\pm 1$ FID on CIFAR‑10).
@@ F.9 Evaluation Metrics
We use FID as the primary quantitative measure of image‑generation quality. FID computes the Fréchet distance between multivariate Gaussian embeddings of real and generated images in the feature space of a pre‑trained Inception‑V3 network (Heusel et al.,, 2017). Lower values indicate that the generated distribution is closer to the real distribution in terms of both perceptual and statistical similarity. Unless noted otherwise, all FID scores are computed on 50k generated samples using the official Inception statistics for each dataset.
@@
Unless stated otherwise, we evaluate every checkpoint with 50k samples and the official Inception statistics. Snapshot selection and the choice of $\lambda_{\mathrm{GP}}$ follow the settings detailed in Sec. S7.
@@ F.10 Logging and Diagnostics
We monitor diagnostics that directly reflect Lipschitz control and training health. In particular, we track the empirical distribution of $\|\nabla_{\tilde{x}}\tilde{h}_{\theta}(\tilde{x})\|_{2}$ on the interpolants used for the penalty, the value of the penalty term itself, and the ratio $\mathbb{E}[(\|\nabla\|-1)^{2}]/\mathbb{E}[\text{critic term}]$. Healthy runs concentrate gradient norms near $1$; persistent mass well above $1$ suggests increasing $\lambda_{\mathrm{GP}}$ or applying the penalty more frequently, whereas mass well below $1$ suggests reducing $\lambda_{\mathrm{GP}}$ or switching to a one‑sided penalty.
@@ F.12 Failure Modes and Stability Notes
The non‑Lipschitz BOLT variant ($\lambda_{\mathrm{GP}}=0$) frequently diverges or produces very high FID, which is consistent with optimizing a TV‑strength objective. Combining strong spectral normalization with a large gradient‑penalty coefficient can over‑regularize the critic; in such cases we prefer to rely primarily on a single Lipschitz control. We also avoid BatchNorm in the critic. If the critic saturates, reducing the discriminator learning rate or increasing $n_{\mathrm{d}}$ usually restores stable training.

### v2

@@ D.5 Other gradient penalties and related methods
  • One‑sided GP (Lipschitz penalty). Penalize only $\|\nabla\tilde{h}_{\theta}\|_{2}>1$: $\lambda\,\mathbb{E}[\max(0,\|\nabla\tilde{h}_{\theta}\|-1)^{2}]$, avoiding a push toward <1<1 gradients that can reduce capacity (Petzka et al., 2018).
  • Zero‑centered penalties $R_{1}/R_{2}$. Apply $\lambda\,\mathbb{E}_{x\sim P_{\mathrm{data}}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{1}$) or $\lambda\,\mathbb{E}_{x\sim P_{g}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{2}$) to improve convergence empirically—even though they do not strictly enforce $1$-Lipschitzness (Mescheder et al., 2018).
  • DRAGAN (local penalties). Sample $\hat{x}=x_{\mathrm{real}}+\delta$ with small noise and penalize $\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}$ near the data manifold to promote local smoothness (Kodali et al., 2017).
  • Orthogonal/Parseval constraints. Constrain $W^{\top}W\approx I$ (Parseval/Björck) to shrink $\|W\|_{2}$ toward $1$ at higher compute cost (Cissé et al., 2017).
  • Lipschitz activations. Using activations with known Lipschitz constants (e.g., GroupSort) can yield provably $1$-Lipschitz networks when combined with spectral control of linear layers (Anil et al., 2019).
@@ Appendix E Additional Experimental Results
E.1 CelebA–64 $\lambda_{\rm{GP}}$ Sensitivity for BOLT–GAN
For completeness, Figure 4 presents the CelebA–64 $\lambda_{\rm{GP}}$ sweep mirroring the CIFAR-10 analysis in the main text. The stability/performance trends are consistent: $\lambda_{\rm{GP}}\!\in\![5,10]$ is best, with variations within $\pm 1$ FID.
@@ E.2 Qualitative Results
We provide qualitative samples to complement the quantitative results in the main paper: each panel below shows an uncurated $4\times 4$ grid of random images generated by BOLT–GAN–GP after 20 training epochs under the same setup as our primary experiments (same architectures, optimizer, and $\lambda_{\rm{GP}}$).
Figure 5: Qualitative samples (first experiment of the main paper). BOLT-GAN-GP after 20 epochs with $\lambda_{\rm{GP}}{=}10$. Top-left: CIFAR-10; top-right: CelebA-64; bottom-left: LSUN Bedroom-64; bottom-right: LSUN Church-64. Each panel shows the same $4\times 4$ grid of generated samples at a reduced scale. Under the identical setup used in our main experiments, Lipschitz BOLT–GAN produces sharper textures, fewer checkerboard artifacts, and more stable color statistics than the WGAN–GP baseline.
@@ F.5 Lipschitz Enforcement
$\mathrm{GP}\;=\;\frac{\lambda_{\rm{GP}}}{B}\sum_{j=1}^{B}\Big(\,\|\nabla_{\tilde{x}_{j}}\tilde{h}_{\theta}(\tilde{x}_{j})\|_{2}-1\Big)^{2}$
to the critic objective. We set $\lambda_{\rm{GP}}=10$ as our default because this value is the de facto choice introduced by WGAN‑GP and widely adopted in subsequent implementations; it offers a robust trade‑off between stability and capacity in practice (Gulrajani et al., 2017; Mescheder et al., 2018). In Appendix F we also report a small sweep over $\lambda_{\rm{GP}}\in\{1,5,10,20\}$ that confirms the default is near‑optimal under our settings.
@@ F.8 Hyperparameter Sweeps and Selection
We sweep $\lambda_{\rm{GP}}\in\{1,5,10,20\}$ separately for each method and dataset, and we report the best FID per method within a fixed compute budget. Other knobs such as $\beta_{1}$ and $n_{\mathrm{d}}$ have a comparatively minor effect in our ranges (e.g., within $\pm 1$ FID on CIFAR‑10).
@@ F.9 Evaluation Metrics
We use FID as the primary quantitative measure of image‑generation quality. FID computes the Fréchet distance between multivariate Gaussian embeddings of real and generated images in the feature space of a pre‑trained Inception‑V3 network (Heusel et al., 2017). Lower values indicate that the generated distribution is closer to the real distribution in terms of both perceptual and statistical similarity. Unless noted otherwise, all FID scores are computed on 10k generated samples using the official Inception statistics for each dataset.
@@
Unless stated otherwise, we evaluate every checkpoint with 10k samples and the official Inception statistics. Snapshot selection and the choice of $\lambda_{\rm{GP}}$ follow the settings detailed in Appendix F.
@@ F.10 Logging and Diagnostics
We monitor diagnostics that directly reflect Lipschitz control and training health. In particular, we track the empirical distribution of $\|\nabla_{\tilde{x}}\tilde{h}_{\theta}(\tilde{x})\|_{2}$ on the interpolants used for the penalty, the value of the penalty term itself, and the ratio $\mathbb{E}[(\|\nabla\|-1)^{2}]/\mathbb{E}[\text{critic term}]$. Healthy runs concentrate gradient norms near $1$; persistent mass well above $1$ suggests increasing $\lambda_{\rm{GP}}$ or applying the penalty more frequently, whereas mass well below $1$ suggests reducing $\lambda_{\rm{GP}}$ or switching to a one‑sided penalty.
@@ F.12 Failure Modes and Stability Notes
The non‑Lipschitz BOLT variant ($\lambda_{\rm{GP}}=0$) frequently diverges or produces very high FID, which is consistent with optimizing a TV‑strength objective. Combining strong spectral normalization with a large gradient‑penalty coefficient can over‑regularize the critic; in such cases we prefer to rely primarily on a single Lipschitz control. We also avoid BatchNorm in the critic. If the critic saturates, reducing the discriminator learning rate or increasing $n_{\mathrm{d}}$ usually restores stable training.